OpenStudy (anonymous):

Simplify (1 + i)(1 - i)

OpenStudy (anonymous):

we know that (a+b)(a-b) =a^2-b^2 hence the answer is 1^2-i^2 =1-(-1) =2

OpenStudy (anonymous):

hey friend @BABYShawol184 you mad a mistake 1 line before Therefore... +i x -i ==i^2

OpenStudy (anonymous):

Thank you @irshadno1 :)

OpenStudy (anonymous):

Well.... (1+i)(1−i) When multiplying binomials we can use F.O.I.L Meaning you multiply the F: first term, the O: outer terms, the I: inner terms and the L: last terms so... (1+i)(1−i) (+1)(+1)+(+1)(−i)+(+i)(+1)+(+i)(−i) Therefore... \[1-i+i-i ^{2}\] Simplified to ... \[1-i^{2}\]